OpenStudy (anonymous):

A 60 kg person stands on a scale in an elevator. How many newtons does the scale read when the elevator is ascending at a rate.of 1 m/s

OpenStudy (nubeer):

Assuming constant acceleration: At rest the scale would read: F=mg, =60kg×9.8ms^2=588N If you are accelerating, regardless of direction, you have a net force acting upon you. Since we are asked about the scale, if you are being accelerated upward then an equal force downward is being applied to the scale, which is also the same direction as the force of gravity: Fnet=ma+mg, (1ms^2×60kg)+(60kg×9.8ms^2)=648N If you are being accelerated downward, again an equal and opposite force is being applied to the scale, the opposite direction of gravity: Fnet=(60kg×(−1ms2))+(60kg×9.8ms2)=528N The last two have a velocity as constant, which means your acceleration is 0: Fnet=ma+mg = 0+60kg×9.8ms2 =588N